OpenStudy (anonymous):

To what volume should you dilute 133mL of an 8.00M CuCl2 solution so that 48.5mL of the diluted solution contains 4.48g CuCl2?

OpenStudy (abmon98):

Dilution Law C1*V1=C2*V2 Number of Moles=Mass/Molar Mass 4.48/134=0.033 moles Number of Moles=Concentration*Volume(dm^3) 48.5/1000*C=0.033 C=0.033*1000/48.5=0.68mol/dm^3 0.68*V2=133/1000*8